OpenStudy (anonymous):

1) How can you use a table to determine the half-life? 2) How can you use a graph to determine the half-life? 3) Which method do you consider most appropriate for determining half-life? 4) If you had a graph of the total sample that had the original sample size at 64 g, and after 72 hours there were only 8 grams left of the substance, what would the half-life of the sample be? Thank you!!! :)

OpenStudy (australopithecus):

Half life is an exponential function, if you know two points in time and what percentage of the substance is left you can determine the function for which half life is determined f(t) = ne^rt where, f(t) = Amount of substance left at time t n = initial amount of substance r = rate t = time

OpenStudy (australopithecus):

sorry I made a mistake in the function it should be f(t) = ne^(-rt)

OpenStudy (australopithecus):

you can figure out half life by using the formula ln(2)/r = halflife

OpenStudy (australopithecus):

using this formula you can extrapolate half life from a graph and a table

OpenStudy (australopithecus):

I hope this is helpful